Determine the nature of the equation: 3x^2-4x-5=0
Answer by ewatrrr(24785)

 
Hi,
3x^2-4x-5=0
y = 3(x-4/6)^2 -16/12 - 60/12
y = 3(x-2/3)^2 - 38/6 Parabola V(2/3, -38/6) line of symmetry x = 2/3